# Einstein metrics on homogeneous superspaces

### v1

@@ 3.2 Riemannian, Ricci and scalar curvatures
Let $g=\langle-,-\rangle$ be a left-invariant graded Riemannian metric on $\operatorname{SL}(1|1)$. Then, $g$ can be identified with a scalar superproduct on $\mathfrak{sl}(1|1)$, which is a non-degenerate supersymmetric even bilinear form. This implies that $\langle X_{1},X_{2}\rangle=\langle X_{1},X_{3}\rangle=\langle X_{2},X_{2}\rangle=\langle X_{3},X_{3}\rangle=0$ and $\langle X_{2},X_{3}\rangle=-\langle X_{3},X_{2}\rangle$, so relative to the ordered basis $\{X_{1},X_{2},X_{3}\}$, we may assume that the metric $g$ has the matrix form
@@ 3.3 Curvature formulas
Let $G=(G_{0},\mathfrak{g})$ be a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra $\mathfrak{g}^{\mathbb{C}}$, and $Q$ a non-degenerate supersymmetric even bilinear form on $\mathfrak{g}^{\mathbb{C}}$, with $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariance given by
@@
Let $K=(K_{0},\mathfrak{k})$ be a closed connected Lie subsupergroup of $G$, and let the homogeneous superspace $M=G/K$ be equipped with a $G$-invariant graded Riemannian metric $\langle$-,$\rangle$ - , - ⟩. Our goal here is to give explicit formulas for the various curvatures of this metric.
We choose once and for all a $Q$-orthogonal complement $\mathfrak{m}$ of $\mathfrak{k}$. We thus have the vector-superspace decomposition $\mathfrak{g}=\mathfrak{k}\,\oplus\,\mathfrak{m}$ and ditto isomorphisms $T_{K}(M)\cong\mathfrak{m}\cong\mathfrak{g}/\mathfrak{k}$ given in terms of Killing fields, as explained below. On $\mathfrak{m}$, the $\operatorname{Ad}_{G}$-representation restricts to the corresponding $\operatorname{Ad}_{K}$-representation, called the isotropy representation.
The transitive action of $G$ on $M$ induces a surjective even linear map from $\mathfrak{g}$ to $T_{K}M$; see (2.4). Correspondingly, we identify the Lie subsuperalgebra $\mathfrak{k}$ of $\mathfrak{g}$ with the Lie superalgebra of those graded Killing vector fields on $M$ which vanish at $K$, and similarly identify nonzero $X\in\mathfrak{m}$ with the nonzero vector field $X^{\dagger}$ for which $(X^{\dagger})_{K}\in T_{K}M$. With [,][-,-][ - , - ] denoting the Lie superbracket of vector fields on $M$, and $[-,-]_{\mathfrak{g}}$ the Lie superbracket of $\mathfrak{g}$, the anti-homomorphicity of (2.3) then implies that
@@
We denote by $Z_{\mathfrak{k}}$, respectively $Z_{\mathfrak{m}}$, the $\mathfrak{k}$- and $\mathfrak{m}$-components of $Z\in\mathfrak{g}$, so
@@
In the following, we primarily use $X,Y,Z,\ldots$ to denote elements of $\mathfrak{g}$, despite the use of the same symbols to represent vector fields in Section 3.1 and Section 3.2. We hope that this practice will not cause any confusion.
The next result gives the Levi-Civita connection on a homogeneous superspace. We follow [Bes87, Proposition 7.28] on classical homogeneous spaces and introduce the even bilinear map $U:\mathfrak{m}\times\mathfrak{m}\to\mathfrak{m}$ given implicitly by
@@
To reduce the number of fractions in the ensuing expressions, especially in Proposition 3.10 and Section 3.4, we have opted to scale $U$ by $2$ relative to the similar function in [Bes87].
@@
Let $M=G/K$ be a homogeneous superspace with a $G$-invariant graded Riemannian metric $\langle$-,$\rangle$ - , - ⟩, and let $X^{\dagger},Y^{\dagger}$ be graded Killing vector fields associated with $X,Y\in\mathfrak{m}$. Then,
@@
$2\,\langle\nabla_{X^{\dagger}}Y^{\dagger},Z^{\dagger}\rangle=-\langle X^{\dagger},[Y,Z]_{\mathfrak{g}}^{\dagger}\rangle-\langle[X,Y]^{\dagger}_{\mathfrak{g}},Z^{\dagger}\rangle-(-1)^{[X^{\dagger}][Y^{\dagger}]}\langle Y^{\dagger},[X,Z]^{\dagger}_{\mathfrak{g}}\rangle.$
@@
By Theorem 3.3, the metric $g$ on $M$ can be identified with a non-degenerate $\operatorname{Ad}_{K}$-invariant supersymmetric even bilinear form on $\mathfrak{m}$. Note that the restriction $g|_{\mathfrak{m}_{\bar{0}}\times\mathfrak{m}_{\bar{0}}}$ may be indefinite and that $\dim(\mathfrak{m}_{\bar{1}})$ is even. Furthermore, by supersymmetry, this bilinear form is symmetric on $\mathfrak{m}_{\bar{0}}\times\mathfrak{m}_{\bar{0}}$ and skew-symmetric on $\mathfrak{m}_{\bar{1}}\times\mathfrak{m}_{\bar{1}}$. As the bilinear form is even, we have $\langle\mathfrak{m}_{\bar{0}},\mathfrak{m}_{\bar{1}}\rangle=\langle\mathfrak{m}_{\bar{1}},\mathfrak{m}_{\bar{0}}\rangle=0$. The $\operatorname{ad}_{\mathfrak{k}}$-invariance of $\langle$-,$\rangle$ - , - ⟩ means that $\langle X,[Y,Z]_{\mathfrak{m}}\rangle=\langle[X,Y]_{\mathfrak{m}},Z\rangle$ for all $X,Z\in\mathfrak{m}$ and $Y\in\mathfrak{k}$.
It is clear that $\mathfrak{m}$ admits an ordered $g$-normalised homogeneous basis $\{X_{1},\dots,X_{c},X_{c+1},\dots,X_{d}\}$, where $[X_{i}]=\bar{0}$ for $i\leq c$ and $[X_{i}]=\bar{1}$ for $i>c$, such that the corresponding matrix representation of $g$ is of the form
@@
where $p+q=c$ and $2r=d-c$. It follows that the right dual basis elements are given by
$\begin{array}[]{rllll}&\widebar{X}_{i}=X_{i},&1\leq i\leq p;&\widebar{X}_{i}=X_{i+r},&c<i\leq c+r,\\[5.69046pt] &\widebar{X}_{i}=-X_{i},&p<i\leq c;&\widebar{X}_{i}=-X_{i-r},&d-r<i\leq d.\end{array}$ (3.14)
@@
Using the super-skew-symmetry of $[-,-]_{\mathfrak{m}}$, it readily follows from Lemma 3.8 that
@@
Using the $\operatorname{ad}_{\mathfrak{k}}$-invariance of $\langle$-,$\rangle$ - , - ⟩, it also follows that, for every $Y\in\mathfrak{k}_{\bar{0}}$,
@@
We are now in a position to state and prove our main curvature formulas: Proposition 3.10, Theorem 3.11 and Corollary 3.12. The proof of the expression for the Ricci curvature given in Theorem 3.11 is lengthy and deferred until Section 3.4.
@@
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded Riemannian metric $\langle$-,$\rangle$ - , - ⟩, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K=(K_{0},\mathfrak{k})$ a connected closed Lie subsupergroup of $G$. Let $\mathfrak{m}$ be a $Q$-orthogonal complement of $\mathfrak{k}$, where $Q$ is an $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant non-degenerate supersymmetric even bilinear form on $\mathfrak{g}^{\mathbb{C}}$, and let $X_{i},X_{j},X_{k},X_{\ell}\in\mathfrak{m}$. Then,
$4\,\langle R$ $(X_{i},X_{j})X_{k},X_{\ell}\rangle=-2\,\langle[X_{i},X_{j}]_{\mathfrak{m}},[X_{k},X_{\ell}]_{\mathfrak{m}}\rangle+(-1)^{[X_{i}]([X_{j}]+[X_{k}])}\langle[X_{j},X_{k}]_{\mathfrak{m}},[X_{i},X_{\ell}]_{\mathfrak{m}}\rangle$
@@
With the shorthand notation $\nabla_{i}=\nabla_{X_{i}}$ and $[i]=[X_{i}]$ for each $i$, and applying (3.1) and (3.4) to the defining expression for the Riemannian curvature, we get
@@
which follow from the $\operatorname{ad}_{\mathfrak{k}}$-invariance. ∎
Although simplifying the ensuing expression for the Ricci curvature is arduous, the formula in Theorem 3.11 below is remarkably similar to its non-super counterpart (cf. [Bes87, Corollary 7.38]).
@@
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded Riemannian metric $\langle$-,$\rangle$ - , - ⟩, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K$ a connected closed Lie subsupergroup of $G$. Let $\{X_{1},\ldots,X_{d}\}$ be a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$. Then, for every $i\in\{1,\ldots,d\}$,
@@
Let the notation be as in Theorem 3.11. Then, the scalar curvature of $\langle$-,$\rangle$ - , - ⟩ satisfies
$S=-\frac{1}{2}\sum_{i=1}^{d}B(\widebar{X}_{i},X_{i})+\frac{1}{4}\sum_{i,j=1}^{d}\langle[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}},[X_{j},X_{i}]_{\mathfrak{m}}\rangle.$
@@ 3.4 Proof of the Ricci curvature formula
This section is devoted to proving the Ricci curvature formula given in Theorem 3.11. Although our general strategy is similar to that in [Bes87, Chapter 7], the super-setting complicates matters considerably. Throughout, $\{X_{1},\ldots,X_{d}\}$ denotes a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$, where $[\widebar{X}_{i}]=[X_{i}]$ for all $i$. As in the proof of Proposition 3.10, we will use the shorthand notation $\nabla_{i}=\nabla_{X_{i}}$ and $[i]=[X_{i}]$. We have used the $Q$-orthogonal decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$ to extend the superproduct $g=\langle-,-\rangle$ on $\mathfrak{m}$ to an $\operatorname{Ad}_{K}$-invariant superproduct on $\mathfrak{g}$ by choosing on $\mathfrak{k}$ an $\operatorname{Ad}_{K}$-invariant superproduct, such as $Q|_{\mathfrak{k}}$, and setting $\langle\mathfrak{k},\mathfrak{m}\rangle=0$. Since the extended superproduct $\langle-,-\rangle:\mathfrak{g}\times\mathfrak{g}\to\mathbb{R}$ is an even bilinear map, an expression like $\langle Y_{1},[Y_{2},Y_{3}]\rangle$, with $Y_{1},Y_{2},Y_{3}\in\mathfrak{g}$, is zero if $[Y_{1}]\neq[Y_{2}]+[Y_{3}]$. We first establish a sequence of technical results: Lemma 3.13 to Lemma 3.17.
@@
[…]
@@
Lemma 3.15.
For each $i=1,\ldots,d$, we have
Proof.
Using (3.10), we have
$\sum_{j=1}^{d}\,\langle U(X_{i},\widebar{X}_{i}),U(\widebar{X}_{j},X_{j})\rangle$ $=-\sum_{j=1}^{d}\Big{(}\langle X_{i},[\widebar{X}_{i},U(\widebar{X}_{j},X_{j})]_{\mathfrak{m}}\rangle+(-1)^{[i]}\langle\widebar{X}_{i},[X_{i},U(\widebar{X}_{j},X_{j})]_{\mathfrak{m}}\rangle\Big{)}=0,$
where the last equality follows from Lemma 3.13. ∎
@@
Use (3.9) and $[X,[Y,\widebar{Y_{k}}]_{\mathfrak{k}}]_{\mathfrak{g}}\in\mathfrak{m}$ to rewrite the sum over $k$ as
@@
where the first equality follows from $[Y,\widebar{Y_{k}}]_{\mathfrak{m}}=\sum_{i=1}^{d}\,\langle X_{i},[Y,\widebar{Y_{k}}]_{\mathfrak{m}}\rangle\widebar{X}_{i}$, and the second from $[X,\widebar{X}_{i}]_{\mathfrak{k}}=\sum_{k=1}^{h}\,\langle Y_{k},[X,\widebar{X}_{i}]_{\mathfrak{k}}\rangle\,\widebar{Y}_{k}$. By Lemma 3.8, this establishes the desired expression. ∎
@@
Since the $g$-normalised basis and its right-dual companion are related as in (3.14), we can use (3.6) to compute $\operatorname{Ric}(X_{i},\widebar{X}_{i})$. Applying Lemma 3.15, Lemma 3.16, Lemma 3.8, (3.9) and (3.15) to the ensuing expression for $\operatorname{Ric}(X_{i},\widebar{X}_{i})$, and using (3.21), yields

### v2

@@ 3.2 Invariant metrics
Let $g$ be a left-invariant graded Riemannian metric on $\operatorname{SL}(1|1)$. Then, $g$ can be identified with a scalar superproduct $\langle-,$-$\rangle$ on $\mathfrak{sl}(1|1)$, which is a non-degenerate supersymmetric even bilinear form. This implies that $\langle X_{1},X_{2}\rangle=\langle X_{1},X_{3}\rangle=\langle X_{2},X_{2}\rangle=\langle X_{3},X_{3}\rangle=0$ and $\langle X_{2},X_{3}\rangle=-\langle X_{3},X_{2}\rangle$, so relative to the ordered basis $\{X_{1},X_{2},X_{3}\}$, we may assume that the metric $g$ has the matrix form
@@ 3.3 Curvature formulas
Let $G=(G_{0},\mathfrak{g})$ be a connected Lie supergroup, with $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra $\mathfrak{g}^{\mathbb{C}}$. As $\mathfrak{g}^{\mathbb{C}}$ is basic classical, there exists a non-degenerate supersymmetric even bilinear form $Q$ on $\mathfrak{g}^{\mathbb{C}}$, with $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariance given by
@@
Let $K=(K_{0},\mathfrak{k})$ be a closed connected Lie subsupergroup of $G$, and let the homogeneous superspace $M=G/K$ be equipped with a $G$-invariant graded Riemannian metric $g$. Our goal here is to give explicit formulas for the various curvatures of this metric at the point $K$ of $M$. As discussed following Theorem 3.6, these formulas determine the full curvature tensors.
Since $M$ is equipped with a $G$-invariant metric, the image under the anti-homomorphism (2.5) of the Lie superalgebra $\mathfrak{g}$ of left-invariant vector fields is contained in the Lie superalgebra of graded Killing fields, cf. Definition 3.5. As $\mathfrak{g}$ is simple, the anti-homomorphism, mapping $X\mapsto X^{\dagger}$, is injective, so we can identity the left-invariant vector field $X\in\mathfrak{g}$ with the Killing vector field $X^{\dagger}$ on $(M,g)$. With [,][-,-] denoting the Lie superbracket of vector fields on $M$, and $_{\mathfrak{g}}$ the Lie superbracket of $\mathfrak{g}$, the anti-homomorphicity then implies that
@@
We now identify the Lie subsuperlagebra $\mathfrak{k}$ of $\mathfrak{g}$ with the subsuperalgebra of the Killing vector fields $X^{\dagger}$ which vanish at the base point $K$, and choose once and for all a $Q$-orthogonal complement $\mathfrak{m}$ of $\mathfrak{k}$. We thus have the vector superspace decomposition $\mathfrak{g}=\mathfrak{k}\,\oplus\,\mathfrak{m}$ and ditto isomorphisms $\mathfrak{g}/\mathfrak{k}\cong\mathfrak{m}\cong T_{K}M$, where we identify $\mathfrak{m}$ with $T_{K}M$ by evaluating the corresponding graded Killing vector field at the point $K$. In this way, the isotropy representation of $K$ in $T_{K}M$ is identified with the restriction of the $\operatorname{Ad}_{G}$-representation of $K$ to $\mathfrak{m}$.
We denote by $\langle-,$-$\rangle$ the scalar superproduct on $\mathfrak{m}$ induced by $g_{K}$ through the identification of $\mathfrak{m}$ with $T_{K}M$. By Theorem 3.6, $\langle-,$-$\rangle$ is an $\mathrm{Ad}_{K}$-invariant, supersymmetric, even bilinear form on $\mathfrak{m}$. As $K$ is connected, the $\mathrm{Ad}_{K}$-invariance is equivalent to $\operatorname{ad}_{\mathfrak{k}}$-invariance, that is,
$\langle X,[Y,Z]_{\mathfrak{m}}\rangle=\langle[X,Y]_{\mathfrak{m}},Z\rangle,$
for all $X,Z\in\mathfrak{m}$ and $Y\in\mathfrak{k}$. We extend the superproduct $\langle-,$-$\rangle$ on $\mathfrak{m}$ to an $\operatorname{Ad}_{K}$-invariant superproduct on $\mathfrak{g}$ by choosing on $\mathfrak{k}$ an $\operatorname{Ad}_{K}$-invariant superproduct, such as $Q|_{\mathfrak{k}}$, and setting $\langle\mathfrak{k},\mathfrak{m}\rangle=0$. In the following, we primarily use $X,Y,Z,\ldots$ to denote elements of $\mathfrak{g}$, despite the use of the same symbols to represent vector fields in Section 3.1 and Section 3.2. We hope this practice will not cause any confusion.
The next result evaluates the Levi-Civita connection on $(M,g)$ at the base point $K$. We denote by $Z_{\mathfrak{k}}$, respectively $Z_{\mathfrak{m}}$, the $\mathfrak{k}$- and $\mathfrak{m}$-components of $Z\in\mathfrak{g}$, so
@@
We follow [Bes87, Proposition 7.28] on classical homogeneous spaces and introduce the even bilinear map $U:\mathfrak{m}\times\mathfrak{m}\to\mathfrak{m}$ given implicitly by
@@
To reduce the number of fractions in the ensuing expressions, especially in Proposition 3.12 and Section 3.4, we have opted to scale $U$ by 22 relative to the similar function in [Bes87].
@@
Let $M=G/K$ be a homogeneous superspace with a $G$-invariant graded Riemannian metric $g$, and let $X^{\dagger},Y^{\dagger}$ be graded Killing vector fields associated with $X,Y\in\mathfrak{m}$. Then,
@@
$2g(\nabla_{X^{\dagger}}Y^{\dagger},Z^{\dagger})=-g(X^{\dagger},[Y,Z]_{\mathfrak{g}}^{\dagger})-g([X,Y]^{\dagger}_{\mathfrak{g}},Z^{\dagger})-(-1)^{[X^{\dagger}][Y^{\dagger}]}g(Y^{\dagger},[X,Z]^{\dagger}_{\mathfrak{g}}).$
@@
We note that the restriction of $\langle-,$-$\rangle$ to $\mathfrak{m}_{\bar{0}}\times\mathfrak{m}_{\bar{0}}$ is symmetric but may be indefinite, that the restriction to $\mathfrak{m}_{\bar{1}}\times\mathfrak{m}_{\bar{1}}$ is skew-symmetric, that $\dim(\mathfrak{m}_{\bar{1}})$ is even, and that $\langle\mathfrak{m}_{\bar{0}},\mathfrak{m}_{\bar{1}}\rangle=\langle\mathfrak{m}_{\bar{1}},\mathfrak{m}_{\bar{0}}\rangle=0$. It follows that $\mathfrak{m}$ admits an ordered $g$-normalised homogeneous basis $\{X_{1},\dots,X_{c},X_{c+1},\dots,X_{d}\}$, where $X_{i}]=\bar{0}$ for $i\leq c$ and $X_{i}]=\bar{1}$ for $i>c$, such that the corresponding matrix representation of $\langle-,$-$\rangle$ is of the form
@@
where $p+q=c$ and 2r=dc2r=d-c. The right dual basis elements are given by
$\begin{array}[]{rllll}&\widebar{X}_{i}=X_{i},\quad&1\leq i\leq p;\qquad&\widebar{X}_{i}=X_{i+r},\quad&c<i\leq c+r,\\[5.69046pt] &\widebar{X}_{i}=-X_{i},\quad&p<i\leq c;\qquad&\widebar{X}_{i}=-X_{i-r},\quad&d-r<i\leq d.\end{array}$ (3.14)
We have the following useful observation.
@@
Using the super-skew-symmetry of $_{\mathfrak{m}}$, it readily follows from Lemma 3.10 that
@@
Using the $\operatorname{ad}_{\mathfrak{k}}$-invariance of $\langle-,$-$\rangle$, it also follows that, for every $Y\in\mathfrak{k}_{\bar{0}}$,
@@
We are now in a position to state and prove our main curvature formulas: Proposition 3.12, Theorem 3.13 and Corollary 3.14. The proof of the expression for the Ricci curvature given in Theorem 3.13 is lengthy and deferred until Section 3.4. We emphasise that the formulas below only give the evaluations of the corresponding curvature tensors at the base point $K$. However, by [San10, Theorem 4.16] (cf. the discussion following Theorem 3.6), this data determines the full curvature tensors.
@@
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded Riemannian metric $g$, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K=(K_{0},\mathfrak{k})$ a connected closed Lie subsupergroup of $G$. Let $\mathfrak{m}$ be a $Q$-orthogonal complement of $\mathfrak{k}$, where $Q$ is an $\operatorname{ad}_{\mathfrak{g}^{\mathbb{C}}}$-invariant non-degenerate supersymmetric even bilinear form on $\mathfrak{g}^{\mathbb{C}}$, and let $X_{i},X_{j},X_{k},X_{\ell}\in\mathfrak{m}$. Then, at the base point $K$,
$4\,\langle(R$ $(X_{i},X_{j})X_{k},X_{\ell}\rangle=-2\,\langle[X_{i},X_{j}]_{\mathfrak{m}},[X_{k},X_{\ell}]_{\mathfrak{m}}\rangle+(-1)^{[X_{i}]([X_{j}]+[X_{k}])}\langle[X_{j},X_{k}]_{\mathfrak{m}},[X_{i},X_{\ell}]_{\mathfrak{m}}\rangle$
@@
With the shorthand notation $\nabla_{i}=\nabla_{X_{i}}$ and $i]=[X_{i}]$ for each $i$, and applying (3.1) and (3.4) to the defining expression for the Riemannian curvature, we get
@@
which follows from the $\operatorname{ad}_{\mathfrak{k}}$-invariance. ∎
@@
Let $M=G/K$ be a homogeneous superspace with $G$-invariant graded Riemannian metric $g$, where $G=(G_{0},\mathfrak{g})$ is a connected Lie supergroup, $\mathfrak{g}$ a compact real form of a basic classical Lie superalgebra, and $K$ a connected closed Lie subsupergroup of $G$. Let $\{X_{1},\ldots,X_{d}\}$ be a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$. Then, at the base point $K$, for every $i\in\{1,\ldots,d\}$,
@@
Let the notation be as in Theorem 3.13. Then, at the base point $K$, the scalar curvature of $g$ satisfies
$S(K)=-\frac{1}{2}\sum_{i=1}^{d}B(\widebar{X}_{i},X_{i})+\frac{1}{4}\sum_{i,j=1}^{d}\langle[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}},[X_{j},X_{i}]_{\mathfrak{m}}\rangle.$
@@ 3.4 Proof of the Ricci curvature formula
This section is devoted to proving the Ricci curvature formula given in Theorem 3.13. Although our general strategy is similar to that in [Bes87, Chapter 7], the super-setting complicates matters considerably. Throughout, $\{X_{1},\ldots,X_{d}\}$ denotes a $g$-normalised homogeneous basis for $\mathfrak{m}=T_{K}(M)$, with right dual basis $\{\widebar{X}_{1},\ldots,\widebar{X}_{d}\}$, where $\widebar{X}_{i}]=[X_{i}]$ for all $i$. As in the proof of Proposition 3.12, we will use the shorthand notation $\nabla_{i}=\nabla_{X_{i}}$ and $i]=[X_{i}]$. Since the extended superproduct $\langle-,-\rangle:\mathfrak{g}\times\mathfrak{g}\to\mathbb{R}$ is an even bilinear map, an expression like $\langle Y_{1},[Y_{2},Y_{3}]\rangle$, with $Y_{1},Y_{2},Y_{3}\in\mathfrak{g}$, is zero if $Y_{1}]\neq[Y_{2}]+[Y_{3}]$. We first establish a sequence of technical results: Lemma 3.15 to Lemma 3.19.
@@
where the last equality uses that $X\in\mathfrak{m}$, $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{m}$ and $\mathfrak{k},\mathfrak{m}]\subseteq\mathfrak{m}$. Since $\mathfrak{g}$ is simple, we have $\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]$, and since $\operatorname{Str}([f_{1},f_{2}])=0$ for all $f_{1},f_{2}\in\mathrm{End}(\mathfrak{g})$, it follows that $\operatorname{Str}_{\mathfrak{g}}(\operatorname{ad}_{\mathfrak{g}}(X))=0$. ∎
Corollary 3.16.
For each $i=1,\ldots,d$, we have
@@
Use (3.9) and $X,[Y,\widebar{Y_{k}}]_{\mathfrak{k}}]_{\mathfrak{g}}\in\mathfrak{m}$ to rewrite the sum over $k$ as
@@
where the first equality follows from $Y,\widebar{Y_{k}}]_{\mathfrak{m}}=\sum_{i=1}^{d}\,\langle X_{i},[Y,\widebar{Y_{k}}]_{\mathfrak{m}}\rangle\widebar{X}_{i}$, and the second from $X,\widebar{X}_{i}]_{\mathfrak{k}}=\sum_{k=1}^{h}\,\langle Y_{k},[X,\widebar{X}_{i}]_{\mathfrak{k}}\rangle\,\widebar{Y}_{k}$. By Lemma 3.10, this establishes the desired expression. ∎
@@
Since the $g$-normalised basis and its right-dual companion are related as in (3.14), we can use (3.6) and Proposition 3.12 to obtain an expression for $\operatorname{Ric}(X_{i},\widebar{X}_{i})$. Applying Lemma 3.16 and (3.15) to the expression for $\mathrm{Ric}(X_{i},\widebar{X}_{i})$, the corresponding terms vanish, yielding
$4\mathrm{Ric}(X_{i},\widebar{X_{i}})$ $=-2\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle[X_{i},X_{j}]_{\mathfrak{m}},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[i]+[j]}\langle[X_{j},\widebar{X}_{i}]_{\mathfrak{m}},[X_{i},\widebar{X}_{j}]_{\mathfrak{m}}\rangle$
$-\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle X_{i},[[X_{j},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X}_{j}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle[X_{i},[X_{j},\widebar{X}_{i}]_{\mathfrak{m}}]_{\mathfrak{m}},\widebar{X}_{j}\rangle$
$+\sum_{j=1}^{d}(-1)^{[j]}\langle X_{j},[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},\widebar{X_{j}}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[j]}\langle[[X_{i},\widebar{X}_{i}]_{\mathfrak{m}},X_{j}]_{\mathfrak{m}},\widebar{X}_{j}\rangle$
$+\sum_{j=1}^{d}(-1)^{[i]+[j]}\langle X_{j},[\widebar{X}_{i},[X_{i},\widebar{X}_{j}]_{\mathfrak{m}}]_{\mathfrak{m}}\rangle+\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle[X_{j},[X_{i},\widebar{X}_{j}]_{\mathfrak{m}}]_{\mathfrak{m}},\widebar{X}_{i}\rangle$
$+2\sum_{j=1}^{d}(-1)^{[j]+[i][j]}\langle X_{i},[X_{j},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle-2\sum_{j=1}^{d}(-1)^{[j]}\langle X_{j},[X_{i},[\widebar{X}_{i},\widebar{X}_{j}]_{\mathfrak{g}}]_{\mathfrak{m}}\rangle$
$+\sum_{j=1}^{d}(-1)^{[i]+[j]}\langle U(X_{j},\widebar{X}_{i}),U(X_{i},\widebar{X}_{j})\rangle.$
Signs may be eliminated by invoking Lemma 3.10 and by applying the skew-supersymmetry of the Lie bracket together with the supersymmetry of the scalar product. Applying Lemma 3.18 and (3.9) to the ensuing expression for $\operatorname{Ric}(X_{i},\widebar{X}_{i})$, and using (3.21), yields